OpenStudy (anonymous):

I'm having a hard time with log and ln and im being given the following problem: log(x+7) + log(32x +4)=3 the problem is asking to solve for x. is there a chance someone can explain to me the steps to solve the problem?

OpenStudy (mathmale):

Rules of logs include the following: log a + log b = log a*b (multiplication) log a - log b = log a/b (division) log a^b = b*log a (exponentiation). Take your log (x+7) + log (32x + 4) and combine them into one log using the first rule listed above.

OpenStudy (cwrw238):

first use laws of logs to make one log log (x+7)(32x + 4) = 3 i'm assuming these are logs to the base 10 so if log = 3 this means its the log of 10^3 so we have (x + 7(32x + 4) = 10^3 = 1000

OpenStudy (mathmale):

Actually, (x+7) * (32x + 4) = 1000. How would you solve that for x? Hint: multiply (x+7)*(32x+4) out first; then set the resulting product = to 1000 and solve that equation for x.

OpenStudy (anonymous):

oh ok thank you guys. it's all very clear now